OpenStudy (anonymous):

where do I begin with proving tan y/(tan^2y-1)=1/(tan y- cot y)

OpenStudy (anonymous):

tany/(tan^y-1) start by dividing both num and denom by tany =1/((tan^y-1)/tany) =1/(tany-(i/tany)) =1/(tany -coty)

OpenStudy (anonymous):

Wow! Thanks. I keep forgetting to split (like in this case) the denominator into to fractions I can work with!